OpenStudy (anonymous):

Rickey has 7 over 8 cup of cereal in a bowl. He then gave two 1 over 4 cups of cereal from his bowl to his little sister. Which of the following best explains the amount of cereal left in Rickey's bowl?

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

4x2=8 so 7/8 -2= 5/8

OpenStudy (anonymous):

So Rickey has \(\large \frac{7}{8}\) and he gives away 2 \(\large \frac{1}{4}\) bowls. This can be shown with the equation \[\large \frac{7}{8}~-~(2 \times \frac{1}{4})=~?\] Can you solve that?

OpenStudy (anonymous):

-2/8 not - 2 sorry

OpenStudy (mathstudent55):

@LegendarySadist You meant: \(\large \dfrac{7}{8}~-~(\dfrac{2}{2} \times \dfrac{1}{4})=\) which then becomes \(\large =\dfrac{7}{8}~-~\dfrac{2}{8} \) \(\large =\dfrac{5}{8} \)

OpenStudy (anonymous):

It says "2 1 over 4 cups". Would that not be \(\large 2 \times \frac{1}{4}\)?

OpenStudy (anonymous):

Since he's giving away 2 \(\large \frac{1}{4} \normalsize th\) cups, that's the same as giving away 1/2 cup. So \[\large \frac{7}{8}~-~\frac{4}{8}~=~\frac{3}{8}\]

OpenStudy (mathstudent55):

Sorry. You are absolutely correct. I didn't notice the two spelled out. I only looked at the numbers.

OpenStudy (mathstudent55):

@WinterBlade1222 Ignore my answer above. Follow @LegendarySadist 's correct answer instead.